• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Makis P, Elena Kh.

    Log entry time 08:05:07 on April10,2007

    Entry number 198461

    keyword=Shift Summary

    People on shift:

    Makis Petratos, SL, till 4:00
    Elena Khrosinkova, SL, after 4:00
    Chiranjib Dutta, TO

    Event:

    02:48 - MCC started to test beam. MCC got up to 60uA, going higher got injector trips.

    03:32 - 04:15 - Harps scans: beam is of order 100micros x and y
    scans posted on elog.

    A few test runs with BeO target.

    04:37 - 04:42 - moved target to He3

    06:11 - beam off: MCC wiil tune it for hall B

    06:12 - 07:20 - changed momentums and right theta, right HVs -> 3.5, right S1 HV -> 74%

    07:20 - MCC has problems to deliver the beam to the hall

    Runs:

    2259 - 2263, Cosmics
    2264 - 2265, BeO tests
    2266 - 2271, prescales rates tests
    2272 - 2280, quasi-elastic